# UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality

### v1

@@ Abstract
Reinforcement Learning from Human Feedback (RLHF) has become a cornerstone for aligning large language models (LLMs) with human values. However, existing approaches struggle to capture the multi-dimensional, distributional nuances of human preferences. Methods such as RiC that directly inject raw reward values into prompts face significant numerical sensitivity issues—for instance, LLMs may fail to distinguish between 9.11 and 9.8—while alternatives like MORLHF, Rewarded Soups, and MODPO incur high computational costs by training multiple models. In this work, we introduce Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel framework that overcomes these limitations. Our approach leverages a diverse set of strictly increasing, non-linear utility functions to transform user-specified preferences into symbolic tokens, which are then used to condition a single LLM. This design not only mitigates numerical reasoning challenges but also substantially reduces training overhead, yielding models that achieve superior Pareto fronts and robust alignment across complex reward dimensions.
UC-MOA: Utility-Conditioned Multi-Objective Alignment for Distributional Pareto-Optimality
Zelei Chengthanks: Equal contribution. Northwestern University Evanston, USA                        Xin-Qiang Cai11footnotemark: 1 RIKEN-AIP Tokyo, Japan                        Yuting Tang The University of Tokyo Tokyo, Japan
Pushi Zhang Microsoft Research Asia Beijing, China                        Boming Yang The University of Tokyo Tokyo, Japan                        Xinyu Xing Northwestern University Evanston, USA
@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF) has emerged as a powerful paradigm for aligning artificial agents with human values by incorporating user-provided feedback into the training process Ouyang et al. (2022); Bai et al. (2022); Stiennon et al. (2020). Despite its success, conventional RLHF approaches typically rely on scalar reward functions and struggle to capture the full spectrum of human preferences, which are inherently multi-dimensional and distributional in nature. For example, methods like RiC (Yang et al., 2024c) directly input raw reward values into the prompts. However, LLMs are known to have limited numerical reasoning capabilities, often failing to distinguish between numerical inputs (e.g., which is bigger? 9.11 versus 9.8 (Yang et al., 2024b)). This numerical insensitivity can lead to suboptimal alignment when dealing with nuanced feedback that spans multiple reward dimensions.
Concurrently, alternative strategies such as MORLHF (Li et al., 2020), Rewarded Soups (Rame et al., 2023), and MODPO (Zhou et al., 2024) attempt to address the complexity of multi-objective alignment by training multiple models to handle different reward combinations Roijers et al. (2013). While these methods provide a richer representation of human preferences, they are computationally expensive and scale poorly with the number of reward dimensions, rendering them impractical for large-scale applications. The high training cost not only limits their real-world deployment but also hinders rapid experimentation and adaptation in dynamic environments.
Figure 1: Illustration of the training processes for existing multi-objective alignment methods compared to our proposed UC-MOA. While methods such as RiC and Rewarded Soups directly inject raw reward values (which are prone to numerical errors) or require multiple models to capture diverse preferences, UC-MOA employs a single model conditioned on a symbolic token derived from a diverse set of strictly increasing utility functions.
To overcome these challenges, we propose Utility-Conditioned Multi-Objective Alignment (UC-MOA) for Distributional Pareto-Optimality. Our framework builds upon recent advances in multi-objective reinforcement learning, where the concept of Pareto-optimality is extended to distributional Pareto-optimality—a notion that prioritizes the entire return distribution over mere expectations Levy (1992); Shaked and Shanthikumar (2007). The difference of our method compared with previous methods are illustrated in Figure 1. Central to our approach is the use of a diverse set of strictly increasing, non-linear utility functions, parameterized via monotonic neural networks Amos et al. (2017); Dugas et al. (2009). These utility functions are trained on offline data with human feedback, enabling them to capture not only the expected values but also higher-order statistical properties such as variance and skewness. By converting user preferences into target rewards through a linear mapping and subsequently conditioning the LLM on a symbolic token (i.e., <max_utility_index> concatenated with the optimal utility index), UC-MOA effectively circumvents the numerical sensitivity issues that plague raw reward injection.
In our framework, the use of utility functions not only resolves the numerical reasoning challenges of previous methods but also introduces a principled mechanism to attain distributional Pareto-optimality. By ranking the outputs of these utility functions based on their percentile scores, we dynamically select the most appropriate utility perspective during training—ensuring that the resulting policy optimally balances conflicting objectives in a distributional sense. This strategy enables a single model to be conditioned on complex, multi-dimensional reward profiles, thereby dramatically reducing the training cost compared to multi-model approaches while achieving superior Pareto fronts. Our empirical evaluations demonstrate that UC-MOA robustly aligns LLM outputs with nuanced human preferences across various tasks, offering a scalable and effective solution for real-world multi-objective alignment.
@@ 2 Related Work
2.1 Large Language Model Alignment
Aligning LLMs with human preferences is a critical challenge, and RLHF has emerged as a cornerstone technique in this domain. Initially popularized by Christiano et al. (2017) using a Proximal Policy Optimization (PPO)-based framework, RLHF has played a key role in developing advanced LLMs such as ChatGPT and Claude (Bai et al., 2022; Ouyang et al., 2022). However, scaling PPO for large models introduces significant efficiency challenges (Choshen et al., 2020; Engstrom et al., 2020), motivating the development of more scalable alternatives. Recent methods like Direct Preference Optimization (DPO) (Rafailov et al., 2023) and the Reward-Ranked Fine-Tuning (RAFT) algorithm (Dong et al., 2023) specifically address alignment challenges in LLMs. DPO directly optimizes preference datasets without relying on explicit reward models, and its online iterative variants further improve exploration and robustness (Xiong et al., 2024; Liu et al., 2024). In parallel, RAFT has been successfully applied in LLMs such as LLaMA3 (Dubey et al., 2024) and Qwen2 (Yang et al., 2024a) by sampling multiple responses per prompt, ranking them via a reward model, and fine-tuning the model on the top-ranked outputs—a strategy shown to enhance reasoning capabilities (Aksitov et al., 2023; Havrilla et al., 2024).
2.2 Multi-objective Alignment
In contrast to traditional RLHF, which optimizes a single reward function, multi-objective alignment aims to balance multiple criteria simultaneously, such as harmlessness, humor, and helpfulness. This approach is crucial for real-world applications where conflicting user preferences and diverse objectives must be integrated into a model’s behavior.
Standard MORLHF (Wang et al., 2024) aggregates multiple reward signals via a linear scalarization strategy, maximizing a weighted sum of scores. However, because this method fixes the weight vector during training, it inherently supports only a single, static user preference. To overcome this limitation and better reflect the nuances of multi-objective alignment, Rewarded Soups (Rame et al., 2023) trains $N$ separate policies—each dedicated to a specific objective—and then linearly interpolates these policies. This process produces a continuous spectrum of models that can be tuned to diverse reward configurations, though it introduces scalability challenges due to the cost of training multiple policies.
Alternatively, RiC Yang et al. (2024c) adapts to different user preferences by conditioning model outputs on multiple reward signals embedded within the prompt. While this method offers flexibility, it still encounters the inherent numerical reasoning limitations of large language models, which struggle to parse and integrate raw numerical reward signals during generation.
Our proposed technique tackles the core challenges of multi-objective alignment by pre-training a diverse set of strictly increasing, non-linear utility functions. We then inject a symbolic token into the prompt that encodes the maximum utility index, enabling the model to indirectly condition its output on the reward structure without relying on precise numerical calculations. This design directly addresses the complexity of multi-objective alignment by allowing the model to seamlessly balance and integrate multiple objectives according to varied user preferences.
2.3 Distributional Reinforcement Learning.
Traditional RL focuses on estimating the expected return of a policy, but distributional reinforcement learning (DRL) extends this framework by modeling the entire distribution of returns rather than just their expectation Bellemare et al. (2017); Zhang et al. (2021). This richer representation of uncertainty has been shown to enhance both learning efficiency and policy quality in various single-objective RL tasks Dabney et al. (2018); Barth-Maron et al. (2018).
Several key algorithms have been developed in this area. Categorical DQN Bellemare et al. (2017) approximates the return distribution using a fixed set of discrete atoms, while Quantile Regression DQN Dabney et al. (2018) models it through quantile regression. Another approach, Distributional MPO Barth-Maron et al. (2018), enhances Maximum a Posteriori Policy Optimization by incorporating distributional learning, leading to improved exploration and stability.
Despite these successes, extending distributional RL to MORL remains an open challenge. The presence of multiple, often conflicting, objectives increases the complexity of reasoning about the full return distribution, as well as how preferences should be expressed over these distributions. Recent work by Cai et al. (2023) introduces Distributional Pareto-Optimal Multi-Objective Reinforcement Learning (DPMORL), which seeks to learn a set of Distributional Pareto-Optimal policies. They show that for any strictly increasing utility function, the policy maximizing expected utility corresponds to a Distributional Pareto-Optimal policy. This result suggests a principled way to incorporate distributional reasoning into MORL while preserving Pareto efficiency. Building on this theoretical foundation, our method distinguishes itself by generating a rich ensemble of non-linear utility functions using monotonic neural networks. By incorporating a symbolic selection mechanism—wherein the maximum utility index is determined and appended to the prompt—we effectively decouple the numerical complexities from the LLM’s generation process. This allows our framework to seamlessly integrate distributional considerations into the multi-objective setting while preserving the desirable properties of Pareto efficiency.
@@ 3 Proposed Technique
In this work, we propose a novel framework that integrates distributional utility modeling with RLHF. Recent work such as RiC (Yang et al., 2024d) and Rewarded Soups (Rame et al., 2023) directly appends raw reward values to the prompts. However, as argued in Section 1, LLMs exhibit limited capability in directly processing precise numerical inputs. Instead, we first train a diverse set of non-linear utility functions from an offline dataset and then introduce a symbolic token into the prompt. This token consists of the index corresponding to the maximum utility value, denoted as <max_utility_index>, concatenated with its associated utility index. During online fine-tuning, after generating a response, we determine the maximal utility index by ranking the outputs of each utility function based on their percentile scores, and then update the LLM via an online process that follows the RiC paradigm except for this novel selection mechanism. Figure 1 shows the framework of our method.
3.1 Utility Function Generation with Human Feedback
The first component of our algorithm is dedicated to constructing a rich set of candidate utility functions that will later guide the derivation of diverse optimal policies. Such diversity is crucial for addressing various distributional preferences and for adapting to different application scenarios. To achieve this, we (i) parameterize each candidate utility function using a monotonic neural network, and (ii) optimize a diversity-promoting loss that encourages discrepancies in both the output values and the local sensitivities of these functions over a normalized domain.
Monotonic Neural Network Parameterization. We denote each candidate utility function by $g_{\psi_{i}}:\mathcal{Z}\rightarrow\mathbb{R}$ for $i=1,\dots,M$, where $\mathcal{Z}\coloneqq[0,1]^{K}$ represents the normalized input space. The network $g_{\psi_{i}}$ is constructed to be monotonic (i.e., non-decreasing) by enforcing non-negativity on its weight matrices and by employing non-decreasing activation functions. This design paradigm follows the approaches proposed in convex neural networks Amos et al. (2017) and monotonic architectures such as DPMORL Cai et al. (2023), which have been shown to approximate any multivariate non-decreasing function with arbitrarily small error Dugas et al. (2009). Implementation details are provided in Appendix B.
@@ 3.1 Utility Function Generation with Human Feedback
Diversity-Promoting Objective. To ensure that the candidate functions span a wide range of utility profiles, we introduce a loss function that encourages dissimilarity both in function values and in their gradients. For a given candidate $g_{\psi_{i}}$, we first define the value discrepancy loss as
@@
where $\mathcal{U}(\mathcal{Z})$ denotes the uniform distribution over $\mathcal{Z}$. Next, to promote differences in the local behavior, we define the gradient discrepancy loss as
$\begin{split}L_{\mathrm{grad}}&(\psi_{i})=\;\min_{j\neq i}\mathbb{E}_{\begin{subarray}{c}\mathbf{z},\mathbf{z}^{\prime}\sim\mathcal{U}(\mathcal{Z})\\ \mathbf{z}\neq\mathbf{z}^{\prime}\end{subarray}}\\ &\left[\frac{g_{\psi_{i}}(\mathbf{z}^{\prime})-g_{\psi_{i}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}-\frac{g_{\psi_{j}}(\mathbf{z}^{\prime})-g_{\psi_{j}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}\right]^{2}.\end{split}$ (2)
@@
where the hyperparameter $\mu\in[0,1]$ balances the relative emphasis on value versus gradient discrepancies. Minimizing $L(\psi_{i})$ via gradient descent for each $i=1,\dots,M$ encourages the ensemble $\{g_{\psi_{i}}\}_{i=1}^{M}$ to cover a broad spectrum of utility profiles over $\mathcal{Z}$.
To satisfy the strict monotonicity condition required for ensuring Distributional Pareto-Optimality, we further modify each candidate by incorporating a small linear term. Specifically, we define the final utility function as
@@
where $\epsilon$ is a small positive constant (e.g., $\epsilon=0.01$ in our experiments). The functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ are then employed as the strictly increasing utility functions for guiding policy optimization.
3.2 Utility Index Selection during Training

### v2

@@ Abstract
Aligning Large Language Models (LLMs) to individual user needs is paramount for effective personalization. However, capturing the diverse and often nuanced preferences of different users presents a significant challenge. Existing approaches may not adequately represent complex preference trade-offs, can be computationally expensive, or rely on the LLM’s limited numerical reasoning capabilities for personalization. We introduce Utility-Conditioned Multi-Objective Alignment (UC-MOA), a novel framework for personalized LLM alignment. UC-MOA develops a diverse set of utility functions, each representing a distinct preference structure, to cater to a wide spectrum of user requirements. These utility functions are constructed with a theoretical guarantee to ensure the resulting set of preference-aligned policies can achieve distributional Pareto-optimality. Instead of conditioning on raw numerical values, UC-MOA employs a “utility-conditioned” method, using a symbolic token that represents an optimal utility function index. This allows the LLM to learn an implicit mapping between the token and the desired personalized response characteristics through fine-tuning, thereby enabling effective personalization. We conduct a user study demonstrating UC-MOA’s superior ability to satisfy diverse user preferences compared to baseline methods, and perform quantitative experiments showing that UC-MOA achieves superior Pareto fronts and maintains high computational efficiency, offering a scalable and robust solution for multi-objective personalized LLM alignment.
@@ 1 Introduction
Large Language Models (LLMs) have demonstrated impressive capabilities across a wide range of natural language tasks and increasingly serve as powerful tools for real-world applications [1, 2, 3]. Developing personalized LLMs are essential for business monetization, as it enables tailored user experiences that enhance engagement, retention, and revenue [4, 5, 6]. However, achieving effective personalization is non-trivial, as it requires addressing diverse and often conflicting user preferences, which cannot be adequately represented by a single scalar reward [7, 8]. Conventional fine-tuning approaches, such as Proximal Policy Optimization (PPO) [9] and Direct Preference Optimization (DPO) [10] optimize a single or aggregated reward signal, which fails to capture the multi-objective nature of real-world user requests. Consequently, there is an urgent need for multi-objective alignment methods that can balance multiple goals to satisfy diverse individual preferences.
Existing methods for multi-objective alignment in LLMs face challenges when applied to fine-grained personalization. Traditional Multi-Objective Reinforcement Learning from Human Feedback (MORLHF) uses linear scalarization of reward signals [11]. This approach struggles to capture complex, non-linear preference structures unique to individual users and typically requires training separate models for different preference weightings, making it computationally prohibitive for wide-scale personalization. More flexible techniques like Rewarded Soups [12] interpolate between multiple fine-tuned models, but still incur the high cost of training and storing multiple LLMs. A recent alternative, Reward-in-Context (RiC) [13] seeks to achieve personalization within a single model by embedding numerical reward targets directly into the prompt, thereby guiding the LLM toward a user-specified point in the reward space. However, a significant drawback of RiC is their reliance on a continuous reward space. To achieve satisfactory personalization performance, the LLM must possess robust numerical reasoning capabilities [14]. Specifically, it needs to accurately judge which training examples or potential generations are quantitatively closer to the user’s multi-dimensional preference vector. This dependence on precise numerical interpretation can be a bottleneck, as LLMs often struggle with such fine-grained quantitative discernment [15], potentially leading to inconsistent or suboptimal personalization outcomes.
To address these limitations and advance personalized LLM alignment, we propose Utility-Conditioned Multi-Objective Alignment (UC-MOA). At the core of UC-MOA is the development of a diverse set of non-linear utility functions, each intended to capture a distinct user preference structure across multi-dimensional reward outcomes. Importantly, this diverse set of utility functions is constructed with theoretical guarantees ensuring that the policies aligned with these utilities can achieve distributional Pareto-optimality [16]. This means the framework can identify policies that offer optimal trade-offs across objectives, considering the entire distribution of outcomes, without being dominated by other achievable policies. Instead of injecting numerical reward targets, which can tax the LLM’s numerical reasoning, we employ a symbolic conditioning mechanism. For each response, we determine which utility function in our diverse set it best satisfies. We then associate this response with a discrete symbolic token, <max_utility_index> $i^{*}$, where $i^{*}$ is the index of that optimal utility function. Through fine-tuning, the LLM learns an implicit mapping between this symbolic token and the nuanced characteristics of responses that align with the corresponding underlying utility function. This strategy enables fine-grained personalization according to diverse preference profiles while sidestepping the LLM’s numerical reasoning limitations.
To summarize, our contributions are:
  • UC-MOA introduces a novel method for personalized LLM alignment by training and leveraging a diverse set of utility functions to simulate varied user preferences. These utility functions are designed to ensure the resulting set of aligned policies can achieve distributional Pareto-optimality.
  • We propose a symbolic utility-conditioned mechanism where the LLM learns a mapping from utility indices to desired response styles, enabling robust personalization while avoiding the pitfalls of direct numerical reward injection.
  • Our comprehensive experiments demonstrate the superiority of UC-MOA in satisfying user preference, achieving competitive Pareto fronts for diverse preference profiles, and maintaining high computational efficiency, offering a scalable and effective solution for personalizing LLMs to complex, multi-faceted user preferences.
@@ 2 Related Work
2.1 Multi-Objective Personalization of LLMs
Traditional LLM alignment, using methods such as PPO [17, 1, 2] and DPO [10], primarily optimizes for a single reward signal. This is ill-suited to the realities of personalized assistants, where users care about several, often conflicting, attributes simultaneously [18]. To address this, MORLHF [11] extends PPO for multiple objectives, often by treating user preferences as different weighting schemes. However, its common reliance on linear scalarization struggles with complex, non-linear preference trade-offs, and scaling to many unique user profiles by training separate models is impractical. Rewarded Soups [12] offer parameter-merging strategies but still face scalability challenges for fine-grained, diverse personalization.
More direct conditioning approaches aim to achieve personalization within a single model. RiC [13] attempts to condition the LLM directly on reward preferences by embedding target reward values into the input prompt. For instance, the model may be instructed to generate responses with specified numerical scores for attributes such as helpfulness and harmlessness. However, the effectiveness of RiC relies heavily on the LLM’s capacity for numerical reasoning, particularly in accurately interpreting continuous input signals, which can present a significant limitation for reliable personalization.
Our UC-MOA framework offers a distinct approach to personalized multi-objective alignment. It begins by learning a library of strictly non-decreasing, non-linear utility functions, each capturing a unique preference structure. The monotonicity constraint ensures that increases in any individual reward dimension do not decrease the overall utility, thereby preserving consistency with rational user behavior. Rather than conditioning the model on numerical reward targets, UC-MOA uses symbolic tokens representing the selected utility functions to guide a single fine-tuned model. This circumvents the LLM’s numerical reasoning limitations while still enabling fine-grained personalization across diverse multi-objective preferences.
2.2 Distributional Pareto-Optimality
Effective multi-objective personalization requires more than optimizing expected reward values, as users often have preferences concerning the distribution of outcomes (e.g., risk, consistency) across objectives [19, 2]. Distributional Pareto-Optimal Multi-Objective RL (DPMORL) [16] provides a formal grounding for distributional Pareto-optimality. DPMORL defines Pareto optimality over return distributions, where a policy is optimal if no other policy stochastically dominates it. Critically, DPMORL shows that policies maximizing any strictly increasing utility function over the multi-objective return distribution are part of the Distributional Pareto set. This links diverse utility functions to the exploration of distributionally optimal trade-offs.
UC-MOA leverages this by constructing a diverse set of utility functions designed to span a comprehensive range of preference structures over distributional outcomes. By conditioning the LLM on symbolic indices tied to these utility functions, UC-MOA implicitly guides the model towards policies that are distributionally Pareto-optimal. This ensures that the achieved personalization is not only diverse but also adheres to a rigorous standard of optimality regarding the full spectrum of possible outcome distributions, without requiring the LLM to perform explicit distributional calculations.
@@ 3 Proposed Technique
Effectively personalizing LLMs to diverse user preferences across multiple objectives is a significant challenge. While recent methods such as RiC [20] attempt personalization by appending raw numerical reward values directly to prompts, LLMs often struggle with precise numerical inputs, limiting their ability to reliably adapt to nuanced user needs, as discussed in Section 1. Our proposed framework, UC-MOA, circumvents this limitation to achieve robust multi-objective personalization.
UC-MOA integrates distributional utility modeling with RLHF. The core idea is to first learn a diverse library of non-linear, strictly increasing utility functions, each representing a distinct personalized preference structure or trade-off. Then, instead of conditioning the LLM on raw numerical values, we use a symbolic token, <max_utility_index>, representing the index of the utility function that best aligns with a desired personalized reward profile. This allows the LLM to learn an implicit mapping between the symbolic token and the characteristics of a personalized response, without requiring explicit numerical reasoning. The online fine-tuning process further refines the LLM’s ability to generate responses consistent with these diverse personalized utility profiles. Figure 1 illustrates the UC-MOA framework.
Figure 1: An illustration of the personalization mechanism employed by UC-MOA. Each user expresses a distinct preference over multiple objectives, such as humor and helpfulness as shown in the figure. These preferences are first mapped to a corresponding utility function from a learned library. The index of the selected utility function is then encoded as a symbolic token and incorporated into the input prompt. This conditioning guides the LLM in generating a response aligned with the target preference profile.
3.1 Library of Non-Decreasing Utility Functions
To enable personalization across a wide spectrum of user needs, the first component of UC-MOA constructs a diverse set of candidate utility functions. These functions play a central role in guiding policy optimization toward various distributional preferences, allowing the LLM to adapt its behavior to different personalized application scenarios and user requirements. We achieve this through two key design choices: (i) parameterizing each candidate utility function using a monotonic neural network, ensuring they can represent complex, non-linear preference structures; and (ii) optimizing a diversity-promoting loss that encourages discrepancies in both the output values and the local sensitivities (gradients) of these functions. Together, these elements ensure that the resulting utility library can capture a broad range of distinct and expressive personalized preference profiles.
Monotonic Neural Network Parameterization. We denote each parameterized utility function by $g_{\psi_{i}}:\mathcal{Z}\rightarrow\mathbb{R}$ for $i=1,\dots,M$, where $\mathcal{Z}\coloneqq[0,1]^{K}$ represents the normalized $K$-dimensional reward space and $\psi_{i}$ denotes the corresponding parameters. The scalar output of $g_{\psi_{i}}(\mathbf{z})$ quantifies the desirability of a reward vector $\mathbf{z}$ under the specific preference structure $i$. Each $g_{\psi_{i}}$ is constructed to be monotonically non-decreasing by enforcing non-negative weights and using non-decreasing activations, following [21, 16]. This allows approximation of any multivariate non-decreasing function, essential for capturing varied personalized preferences. Details are in Appendix D.
@@ 3.1 Library of Non-Decreasing Utility Functions
Diversity-Promoting Objective. To ensure the library ${g_{\psi_{i}}}$ spans a wide range of personalized utility profiles, we optimize a diversity loss $L(\psi_{i})$ that encourages dissimilarity both in function values and in their gradients. We first define the value discrepancy loss as
@@
where $\mathcal{U}(\mathcal{Z})$ denotes the uniform distribution over $\mathcal{Z}$ and $\psi_{i}$ denotes the parameters of $g_{\psi_{i}}$. Next, to promote differences in the local behavior, we define the gradient discrepancy loss as
$\begin{split}L_{\mathrm{grad}}(\psi_{i})=\;\min_{j\neq i}\mathbb{E}_{\begin{subarray}{c}\mathbf{z},\mathbf{z}^{\prime}\sim\mathcal{U}(\mathcal{Z})\\ \mathbf{z}\neq\mathbf{z}^{\prime}\end{subarray}}\left[\frac{g_{\psi_{i}}(\mathbf{z}^{\prime})-g_{\psi_{i}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}-\frac{g_{\psi_{j}}(\mathbf{z}^{\prime})-g_{\psi_{j}}(\mathbf{z})}{\|\mathbf{z}^{\prime}-\mathbf{z}\|}\right]^{2}.\end{split}$ (2)
@@
This loss encourages dissimilarity in both function values and gradients, ensuring distinct preference representations. The hyperparameter $\mu\in[0,1]$ balances these terms.
To ensure robust personalized solutions, we enforce the strict monotonicity required for achieving Distributional Pareto-Optimality, and define the final utility functions as
@@
where $\epsilon$ is a small positive constant, set to $0.01$ in our experiments. The functions $\{\hat{g}_{\psi_{i}}\}_{i=1}^{M}$ constitute a library of strictly increasing utility functions, each encoding a distinct way to value multi-objective outcomes, which is critical for enabling diverse personalization.
3.2 Utility Index Selection via Percentile Ranking